OpenStudy (anonymous):

Can you factor 3x^2-16x+21

OpenStudy (anonymous):

x=3, 7/3

OpenStudy (anonymous):

using the quadratic formula \[(16+(\sqrt{256-4(3)(21)}))/2(3) = 18/6\] and \[(16-(\sqrt{256-4(3)(21)}))/2(3) = 14/6 = 7/3\]